OpenStudy (anonymous):

Which of the following is not equivalent to the formula for the area of a rectangle A = LW? A/WL = 1 A/W = L A = L/W A/L = W

OpenStudy (diyadiya):

A = L/W

OpenStudy (anonymous):

A=L/W you can get A/WL=1 by dividing by WL A/W=L by dividing by L and A/L=w by dividing by L

OpenStudy (anonymous):

just answered this same question here: http://openstudy.com/#/updates/4eef951ee4b0367162f62df7